TERAHERTZ SPECROSCOPY OF METHYLAMINE

\underline{R. A. MOTIYENKO}, L. MARGULÈS, Laboratoire PhLAM, UMR 8523 CNRS - Université Lille 1, 59655 Villeneuve d’Ascq Cedex, France; V. V. ILYUSHIN, E. A. ALEKSEEV, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; B. DROUIN, S. YU, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109, USA; J. CERNICHARO, B. TERCERO, Centro de Astrobiología (CSIC-INTA). Laboratory of Molecular Astrophysics. Department of Astrophysics. Ctra de Ajalvir, Km 4, 28850 Torrejòn de Ardoz, Madrid, Spain.

Methylamine (CH_{3}NH_{2}) is a light molecule with an intense rotational spectrum that extends far beyond 1 THz even at temperatures characteristic for interstellar medium. This fact makes methylamine an interesting object of studies for THz radio telescopes like Herschel and SOFIA. In this context we present recent advances in global fitting of the rotational spectrum of parent isotopic species of CH_{3}NH_{2} up to 2.6 THz. The Hamiltonian used in the analysis is based on the group-theoretical formalism developed by Ohashi and Hougen^{a}. It has been successfully applied for the analysis of mmwave and far-IR spectrum of methylamine^{b} as well as for the analysis of sub-THz spectra of its ^{13}C isotopologue^{c}. We will also present the latest results on searches for interstellar ^{13}CH_{3}NH_{2} as well as the results on the analysis of deuterated species of methylamine which is in progress now. This work is supported by Centre Nationale d’Etudes Spatiales (CNES), Action sur Projet Physico-Chimie du Milieu Interstellaire (PCMI-CNRS) and by the contract ANR-08-BLAN-0054.

aOhashi, N. and Hougen, J. T. J. Mol. Spec. 121 (1987) 474.

bIlyushin, V.V. et al. J. Mol. Spec. 229 (2005) 170.

cR.A. Motiyenko et al. FA08, 66th International Symposium on Molecular Spectroscopy (2011).